OpenStudy (anonymous):

Find the limit for lim x--->4 (4x-x^2) / (2-sqrt x)

OpenStudy (abb0t):

Take the conjugate of the function. Can you do that?

OpenStudy (abb0t):

multiply by conjugate*

OpenStudy (anonymous):

It's 2+sqrt x right?

OpenStudy (abb0t):

Yes.

OpenStudy (jdoe0001):

\(\bf \cfrac{4x-x^2}{2-\sqrt{x}} \times \cfrac{2+\sqrt{x}}{2+\sqrt{x}} \implies \cfrac{x(4-x)(2+\sqrt{x})}{\color{blue}{(2-\sqrt{x})(2+\sqrt{x})}}\\ \textit{keeping in mind that }\color{blue}{(a-b)(a+b)} = (a^2-b^2)\)

OpenStudy (anonymous):

Thanks guys, I think I got it